# SLAC: Simulation-Pretrained Latent Action Space for Whole-Body Real-World RL

### v3

@@ Abstract
Keywords: Real-world RL, Latent Action, Whole-body Manipulation
@@ 2 Related Work
Robots can be trained to perform tasks using four main approaches: (1) sim-to-real reinforcement learning, (2) real-world reinforcement learning, (3) classical motion planning and control, and (4) learning from demonstrations. In this section, we review the first two approaches, which are most relevant to SLAC. We defer discussion of the latter two [2, 3, 1, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] to the appendix.
@@ 3.1 Learning a Latent Action Space in Simulation
where is a set of state entities (e.g. whiteboard, table, body parts) in the environment that the robot can interact with; is the latent action space, factorized by design into dimensions; and is a weighting factor for the disentanglement objective. Intuitively, this objective encourages each latent action dimension to control only its corresponding state entity , thereby creating a disentangled and temporarily extended action space that allows the robot to independently and simultaneously control different entities in the environment – an ability critical for downstream learning.
To optimize this objective tractably, we can approximate Eq. 3 through variational inference [54], resulting in the following reward function:
@@
3.2 Sample-Efficient Policy Learning of Downstream Tasks in the Real World
@@
Factored Q-Function Decomposition: Challenging robotics tasks often come with a naturally composite reward function, where the eventual reward is the sum of a set of reward terms corresponding to a set of sub-objectives, e.g. a whole-body mobile manipulation task may require: (1) navigating to a location, (2) without colliding with obstacles, and (3) while holding an object at the right orientation. Directly optimizing this complex reward function with vanilla RL can be quite challenging, often requiring many steps of environment interactions, posing a significant challenge to real-world learning. Our key realization is that we can take advantage of the disentangled nature of the learned latent action space , by finding and exploiting the strong dependencies between the latent action dimensions and the reward terms. Specifically, given a composite reward function with terms444It is important to note that SLAC is still valid with a non-composite reward (i.e. ), just that we will no longer benefit from Q decomposition.: , it can be shown (proof in the appendix) that:
where each factored value function represents the expected return for a specific reward term .
Now, since each dimension of our latent action is trained to control and only control one environment entity, each reward term typically only depends on a small subset of the latent action dimensions (for example, a reward for navigation is only associated with the latent action dimension that controls the robot base location). This property allows us to dissociate each with unrelated latent action dimensions, resulting in , where is the index set corresponding to the dimensions of that reward depends on.
Such a factorization brings us two significant advantages: First, it effectively prevents from learning spurious correlations with actions, thereby providing a better optimization landscape for accurate prediction of the Q value. Second, since we calculate the policy update by backpropagating from the Q functions, each action dimension will only be updated with reward terms that it can affect, resulting in a more accurate policy gradient estimation than the non-factorized version. In other words, our technique effectively decomposes a hard learning problem into multiple simpler problems that can be solved in parallel, leading to improved performance and sample efficiency.
In practice, we implement the Q-decomposition by masking out irrelevant action dimensions for each using a binary adjacency matrix that encodes action-reward dependencies. This implementation enables parallel computation of all Q-functions, significantly accelerating training. The matrix can either be learned automatically from a small number of random interaction trajectories [10], or manually specified when the mapping is known a priori [13]. In SLAC, we assume ground-truth access to this mapping, leveraging the fact that each latent action dimension is intentionally constructed to control a specific environment entity (see Sec. 3.1).
@@ 5 Conclusion
Limitations: Despite its strong empirical performance, SLAC is not without limitations. First, SLAC introduces an implicit trade-off related to the granularity of the latent action space. For example, an identity mapping between the robot’s raw action space and latent action space would make the downstream task policy capable of learning any task within the original capability of the robot, but would significantly reduce the sample efficiency (as shown in our results in Sec. 4.1). Similarly, the temporal length of each latent action entails another tradeoff, where shorter latent actions give more control to the task policy at the expense of longer task horizons, which may hamper learning (as shown in our ablation studies in Sec. 4.2). It is likely that the optimality of the latent action space will be strongly task-dependent. Second, in the current SLAC framework, the latent action decoder is kept fixed during downstream learning. However, for more fine-grained tasks, we might benefit from finetuning the latent action decoder while training the task policy. While this is conceptually doable via the option framework [64], we leave the empirical study of how to implement it in a stable manner to future work. Finally, while this paper primarily focuses on the algorithmic side of real-world learning, we expect future engineering efforts in the automation of task reset and downstream reward generation (e.g., via VLM/LLM or a learned reward function) to further boost the downstream learning efficiency.
@@ A.1 Policy and Training Visualizations
We encourage the reviewer to visit our project website (https://robo-rl.github.io/) for full videos of the SLAC training process in the real world and the learned policies.
@@ A.2 Properties of the Learned Latent Action Space
A.3 Q Decomposition Proof
@@
A.4 Universal Safety Reward
@@
A.5 SLAC Unsupervised Latent Action Space Learning Pseudocode
@@
A.6 SLAC Real-World Downstream RL Pseudocode
@@ A.10 Discussion of Broader Areas Related to SLAC
Classical Motion Planning and Control: A traditional way to enable robots to perform tasks is through motion planning and control. In practice, however, uncertainty and inaccuracy in localization frequently impede the accurate execution of planned trajectories [45, 46, 47, 48, 49]. Moreover, when the robot needs to consider multiple objectives, creating a motion planner is even harder, as it requires solving complex multi-objective optimization problems [42, 43, 44]. On the side of control, existing methods [34, 35, 36, 37, 38, 39, 40, 41] resort to sophisticated prioritized solutions that require extensive tuning and pre-determined task priorities. Moreover, these methods assume accurate models of the robot and the environment, which often break in unstructured environments and with high-dimensional sensor signals (e.g. images). By comparison, SLAC can autonomously learn closed-loop policies only based on onboard sensors, and does not require prior domain knowledge.

### v4

@@ Abstract
Keywords: Real-world RL, Sim2Real, Whole-body Manipulation
@@ 2 Related Work
Robots can be trained to perform tasks using four main approaches: (1) sim-to-real reinforcement learning, (2) real-world reinforcement learning, (3) classical motion planning and control, and (4) learning from demonstrations. In this section, we review the first two approaches, which are most relevant to SLAC. We defer discussion of the latter two [2, 3, 1, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] to the appendix.
@@ 3.1 Learning a Latent Action Space in Simulation
where is a set of state entities (e.g. whiteboard, table, body parts) in the environment that the robot can interact with; is the latent action space, factorized by design into dimensions; and is a weighting factor for the disentanglement objective. Intuitively, this objective encourages each latent action dimension to control only its corresponding state entity , thereby creating a disentangled and temporarily extended action space that allows the robot to independently and simultaneously control different entities in the environment – an ability critical for downstream learning. To optimize this objective tractably, we can approximate Eq. 3 through variational inference [53], resulting in the following reward function:
@@
3.2 Sample-Efficient Learning of Downstream Tasks in the Real World
@@
Factored Q-Function Decomposition: Challenging robotics tasks often come with a naturally composite reward function, where the eventual reward is the sum of a set of reward terms corresponding to a set of sub-objectives, e.g. a whole-body mobile manipulation task may require: (1) navigating to a location, (2) without colliding with obstacles, and (3) while holding an object at the right orientation. Directly optimizing this complex reward function with vanilla RL can be quite challenging, often requiring many steps of environment interactions, posing a significant challenge to real-world learning. Our key realization is that we can take advantage of the disentangled nature of the learned latent action space , by finding and exploiting the strong dependencies between the latent action dimensions and the reward terms. Specifically, given a composite reward function with terms444Note that SLAC is still valid with a non-composite reward (i.e. ) and without Q decomposition.: , it can be shown (proof in the appendix) that , where each factored value function represents the expected return for a specific reward term . Now, since each dimension of our latent action is trained to control and only control one environment entity, each reward term typically only depends on a small subset of the latent action dimensions (for example, a reward for navigation is only associated with the latent action dimension that controls the robot base location). This property allows us to dissociate each with unrelated latent action dimensions, resulting in , where is the index set corresponding to the dimensions of that reward depends on.
Such a factorization brings us two significant advantages: First, it effectively prevents from learning spurious correlations with actions, thereby providing a better optimization landscape for accurate prediction of the Q value. Second, since we calculate the policy update by backpropagating from the Q functions, each action dimension will only be updated with reward terms that it can affect, resulting in a more accurate policy gradient estimation than the non-factorized version. In other words, our technique effectively decomposes a hard learning problem into multiple simpler problems that can be solved in parallel, leading to improved performance and sample efficiency. In practice, we implement the Q-decomposition by masking out irrelevant action dimensions for each using a binary adjacency matrix that encodes action-reward dependencies. This implementation enables parallel computation of all Q-functions, significantly accelerating training. The matrix can either be learned automatically from a small number of random interaction trajectories [10], or manually specified when the mapping is known a priori [13]. In SLAC, we assume ground-truth access to this mapping, leveraging the fact that each latent action dimension is intentionally constructed to control a specific environment entity (see Sec. 3.1).
@@ 5 Conclusion
6 Limitations and Future Works
Despite its strong empirical performance, SLAC is not without limitations. First, SLAC introduces an implicit trade-off related to the granularity of the latent action space. For example, an identity mapping between the robot’s raw action space and latent action space would make the downstream task policy capable of learning any task within the original capability of the robot, but would significantly reduce the sample efficiency (as shown in our results in Sec. 4.1). Similarly, the temporal length of each latent action entails another tradeoff, where shorter latent actions give more control to the task policy at the expense of longer task horizons, which may hamper learning (as shown in our ablation studies in Sec. 4.2). It is likely that the optimality of the latent action space will be strongly task-dependent. Second, in the current SLAC framework, the latent action decoder is kept fixed during downstream learning. However, for more fine-grained tasks, we might benefit from finetuning the latent action decoder while training the task policy. While this is conceptually doable via the option framework [63], we leave the empirical study of how to implement it in a stable manner to future work. Finally, while this paper primarily focuses on the algorithmic side of real-world learning, we expect future engineering efforts in the automation of task reset and downstream reward generation (e.g., via VLM/LLM or a learned reward function) to further boost the downstream learning efficiency.
Acknowledgments
We thank members of RobIn and LARG for their valuable feedback on real robot setup (Shivin Dass, Arpit Bahety, Rutav Shah) and the manuscript (Shivin Dass, Chris Huang, Yifeng Zhu). This work is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (W911NF-25-1-0065), ARO (W911NF-23-2-0004), DARPA (Cooperative Agreement HR00112520004 on Ad Hoc Teamwork) Lockheed Martin, UT Austin’s Good Systems grand challenge, and funding from Two Sigma Investments, LP. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Two Sigma Investments, LP. Peter Stone serves as the Chief Scientist of Sony AI and receives financial compensation for that role. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
@@ A.1 Policy and Training Visualizations
We encourage the reader to visit our paper website for videos of the SLAC training process in the real world and the learned policies.
@@ A.2 Properties of the Learned Latent Action Space
A.3 Discrete vs Continuous Latent Actions
A.4 Q Decomposition Proof
@@
A.5 Universal Safety Reward
@@
A.6 SLAC Unsupervised Latent Action Space Learning Pseudocode
@@
A.7 SLAC Real-World Downstream RL Pseudocode
@@ A.10 Discussion of Broader Areas Related to SLAC
Classical Motion Planning and Control: A traditional way to enable robots to perform tasks is through motion planning and control. In practice, however, uncertainty and inaccuracy in localization frequently impede the accurate execution of planned trajectories [44, 45, 46, 47, 48]. Moreover, when the robot needs to consider multiple objectives, creating a motion planner is even harder, as it requires solving complex multi-objective optimization problems [41, 42, 43]. On the side of control, existing methods [34, 35, 36, 37, 38, 39, 40] resort to sophisticated prioritized solutions that require extensive tuning and pre-determined task priorities. Moreover, these methods assume accurate models of the robot and the environment, which often break in unstructured environments and with high-dimensional sensor signals (e.g. images). By comparison, SLAC can autonomously learn closed-loop policies only based on onboard sensors, and does not require prior domain knowledge.